Hello!!
So i tried this equation many times but it did not work out. The Question was:
Tyler rows 10km downstream in 2 hours. On the return trip, it takes him 4 hours to travel 8 km. Determine his average rowing speed and the speed of the current.
This is what i tried:
Let x rep. his average rowing speed
let y rep. his speed of current
1 equation: 10=(x=y)2
2nd equation: 8+(x=y)4
but i got it wrong, so I was hoping you guys could help me out. Thank you very much!

Answer by stanbon(75887)

Tyler rows 10km downstream in 2 hours. On the return trip, it takes him 4 hours to travel 8 km. Determine his average rowing speed and the speed of the current.
This is what i tried:
Let x rep. his average rowing speed
let y rep. his speed of current
downstream equation: 10 =(x+y)2
upstream equation: 8 =(x-y)4
-------
x + y = 5
x - y = 2
------
Add and solve for "x":
2x = 7
x = 3.5 km/h (average rowing speed
---
Solve for "y"::
x + y = 5
3.5 + y = 5
y = 1.5 km/h (speed of current)
------------------------
Cheers,
Stan H.
-----------------